OpenStudy (anonymous):

The set of integers {3, 4, 5} is a Pythagorean triple. another such set is?

OpenStudy (anonymous):

you can multiply each number by any fixed constant if you want another one, say \(\{6,8,10\}\)

OpenStudy (anonymous):

but if you want to be more creative and find more that are not multiples of the one you have, there are infinitely many for example \(\{5,12,13\}\)

OpenStudy (anonymous):

okay well this question is multiple choice and the answers I have are 6,7,8 6,8,12 6,12,13 and 8,15,17

OpenStudy (anonymous):

in fact if you want say 7 to be one side, you can write \[m+n=7,m-n=1\implies m=4,n=3\] and then use \[\{7, 2\times 4\times 3,4^2+3^2\}\] giving the triple \[\{7,24,25\}\]

OpenStudy (anonymous):

in that case, your job is not to find one, but to check if \[a^2+b^2=c^2\]

OpenStudy (anonymous):

for example \[8^2+15^2=17^2\] \[64+225=289\] which is true the others will be false

OpenStudy (anonymous):

ohh okay that makes sense so I just have to check it the pythagorean theorem works

OpenStudy (anonymous):

yeah, right

OpenStudy (anonymous):

so the answer would be the {8,15,17}

OpenStudy (anonymous):

yes, obviously.

OpenStudy (anonymous):

thanks so much !

OpenStudy (anonymous):

i have another question

OpenStudy (anonymous):

close this question and post it